# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v3

@@ 3.2 Processing Times
Figure LABEL:fig:speed shows the normalized processing times for the different algorithms and corpus types; a table showing the absolute processing times recorded for all classification procedures implemented can be found in appendix C. To improve plot readability, vertical line segments connecting all processing times obtained by each algorithm for the various corpora have been added. Fastest and slowest processing times for each corpus type are shown in seconds at both extremities of the x-axis. Finally, in order to better highlight the differences between both slower and faster algorithms, the x-axis has been rescaled using arcsine transformation.
@@
In the context of the present project, such differences in performance makes processing speed a crucial dimension of linguistic classification performance. Indeed, while a difference of a few seconds at the sample level may not seem so crucial at first glance, the sheer size of OpenAlex implies that processing speed differences beyond one order of magnitude might likely result in delays ranging from weeks to even months.
Regarding the performances of FastSpell and FastText in particular, the former is slightly slower than the latter in all cases, which can be expected given that FastSpell adds an upstream spelling checking functionality to FastText. As shown in the previous figure, this enhancement translates into a gain in recall performance for languages included in the Other category. The question remains, however, whether these local gains in performance are worth the cost in terms of increased processing times.
As for DetectLanguage, the fact that the algorithm can only be used through API queries has an undeniable impact on the processing speed of all procedures using this algorithm: even though the algorithm’s batch mode helps reduce the number of queries by processing multiple inputs at a time, the fact that its processing speed ultimately depends on factors that have nothing to do with algorithmic performance /textitstricto sensu, for example connection speed or server response time, unfortunately makes it an unviable language detection strategy for the OpenAlex database.
To a lesser extent, differences in processing times can also be observed between the various corpus types. Across all algorithms, the Titles (T) and Titles & Abstracts (A) corpora are the corpus types whose processing were the fastest and slowest respectively, with times of $15.98$ and $20.64$ seconds. Average document size can probably and partly explain such difference: since titles tend to be substantially shorter than abstracts, with $72.26$ characters on average to $1238.85$ characters for the latter, corpora that include abstracts will take more time to process than those that do not. From a performance evaluation perspective, a situation similar the one observed for FastSpell and FastText thus arises: while abstracts can be very informative from a language detection perspective, taking them into account for that purpose is however bound to slow down algorithms; whether this gain in information results in an overall gain of performance will ultimately depend on its impact on precision and recall scores, which once again proves the crucial relevance of corpus processing times to the present language detection algorithm assessment.
@@ 3.3 Simulation Results
Looking first at the topmost subfigure, which shows $F_{\beta,\gamma}$ scores for $\gamma=0$ (where is not taken into account), best-performing procedures for each algorithm show interesting similarities. First and foremost, performance for all algorithms seem to follow a dual performance regime, with a transition point at around $\beta\approx 1$, in other words where the weights of precision and recall converge. In the case of DetectLanguage and LangID, this cut-off point occurs a bit earlier ($\beta\approx 0.6$) and later ($\beta\approx 1.2$) along the $\beta$ continuum, but the regime shift is just as sudden and lasting as elsewhere. The first characteristic of that shift is that that all algorithms switch corpus types: whereas the best-performing procedure of CLD3 transitions from the Titles & Abstracts to the Greedy corpus (i.e. it adds journal names), procedures for all other algorithms become title-based as recall weight increases. The second salient feature of the dual-regime observed at $\gamma=0$ is that, again with the exception of the exponential decay of CLD3, which seems unaffected by recall weight increase, all best-performing procedures switch to an exponentially-decreasing growth regime as they hit their respective cut-off points and transitions to the titles-based corpus.
@@
In light of these considerations and the simulation results from which they are derived, which classification procedure performs best depends on the importance given to the different evaluation measures implemented. As regards to corpus types, emphasizing precision over recall will favor the more metadata-rich procedures (i.e. the greedy and titles & abstracts corpora); however, as recall gains in importance and as soon as processing speed is taken into account, the Titles corpus becomes the unanimous choice. As regards to algorithms, while LangID stands out in precision-oriented, speed-overlooking settings, FastSpell and FastText outperforms other algorithms in all other contexts, and the extent of their predominance increases proportionally with both $\beta$ and $\gamma$.
@@ 4 Discussion
The present project aimed to systematically compare various language classification procedures, procedures combining various Python language detection algorithms and metadata-based corpora extracted from manually-annotated articles sampled from the OpenAlex database. Following an analysis of precision and recall performance for each algorithm, corpus, and language as well as of processing speeds recorded for each algorithm and corpus type, overall procedure performance at the database level was simulated using probabilistic confusion matrices for each algorithm, corpus, and language as well as a probabilistic model of relative article language frequencies for the whole OpenAlex database. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on the greedy corpus gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, the procedure combining the FastSpell algorithm and the Titles corpus outperforms all other alternatives.
@@
In addition to the quality of its overall performance, FastText has the advantage of returning as output the probability value of the most probable language. In this sense, a two- or multi-step classification procedure would be quite feasible. For instance, FastText could be run on the titles of all articles indexed in the database, then other procedures that perform better in specific contexts could be run on all cases where the probability/confidence score of the detected language is below a certain threshold. Given the imbalanced nature of the OpenAlex database and the inherent algorithmic tendency to overpredict the majority class, a special effort could then be made to improve prediction for English documents: as mentioned in cespedes2025evaluating, English documents are overrepresented in the OpenAlex database at a rate of around 9%, which negatively affects the representativity of all other languages in the database to a greater or lesser extent, but also reinforces their underrepresentation in the scholarly and knowledge ecosystem at large. In that perspective and given OpenAlex’s status as the first truly multilingual database, improving the database’s language metadata has social and political implications that go well beyond the functionalities that are normally expected of bibliographic databases.
4.1 Limitations of the study
@@
Results presented and discussed in the previous paragraphs are far from definitive. For once, they are inextricably tied to the current state-of-the-art in the field of automatic language detection. The conclusions drawn are in no way meant to solve the problem of the linguistic classification of OpenAlex articles once and for all, but only to attempt to offer the best possible algorithmic recommendation based on current domain knowledge as well as on the available techniques deemed most suitable for that specific purpose.
@@
Another important point to mention is that since the observations and analyses presented here are the direct result of the methodological choices adopted by the authors, the conclusions derived from the results of this project are valid only within the framework of this precise modus operandi. In other words, different choices as regards to the selection of performance measures or the design and parametrization of the simulation process may well have led to different conclusions. The case of the aggregation of performance scores performed here is suggestive in this regard, as increased differences in weighing between precision, recall, and processing speed could lead to different results; the question remains however as to whether more diversified or extreme weighing regimes than those presented here would be really useful. Nevertheless, a certain epistemic caution is in order here, as the results presented here are far from definitive and should not be considered as such.
@@
As regards to OpenAlex itself, it is also important to note that a significant part of the data contained in the database is provided by the journal publishers themselves based on what they consider relevant bibliometric data. Such decisional idiosyncrasies can have an impact on the performance of monolingual language detection algorithms and the language classification they provide, for example if multilingual titles or abstracts are provided for specific journals or articles. This specific scenario in turn points to an important consideration regarding the functionalities of the algorithms considered here, but also those of language detection algorithms in general. While current monolingual language detection algorithms offer above human-level performance over hundreds of different languages, a current and major challenge in Natural Language Processing consists in designing algorithms that achieve high accuracy when dealing with documents written in multiple languages. Given these considerations, multilingual record entries probably represents the biggest current linguistic classification issue for the OpenAlex database: not only is the quality of linguistic metadata inevitably affected by this problem, but its extent is hard to assess, given the difficulty of identifying which attributes are multilingual and as such likely to cause misclassifications.
@@
Despite these limitations, the results presented here show that, with the right language detection procedure, a high-quality linguistic classification of all articles indexed in OpAl is nevertheless possible, classification which would not only reinforce the latter’s status as the first truly multilingual, large-scale bibliometric database, but also and more generally foster cross-linguistic bibliometric assessments and analyses of unmatched accuracy and comprehensiveness.
@@ Appendix A OpenAlex linguistic distribution, grouped by metadata availability
Language Title
Title &
Abstract
Title &
Journal Name
Title, Abstract &
Journal Name
@@ Abstract
Appendix B Pseudocode description for corpus building procedures
The following table shows the name, symbol, and procedure used to build the different corpora used for this project.
Name Symbol Procedure
Title T
FOREACH article, GET title.
Title & abstract,
else title
TAeT
FOREACH article, GET (title, abstract);
IF NO abstract, GET title.
Abstract, else title AeT
FOREACH article, GET abstract;
IF NO abstract, GET title.
Title & journal name,
else title
TJeT
FOREACH article, GET (title, journal name);
IF NO journal name, GET title.
Journal name,
else title
JeT
FOREACH article, GET journal name;
IF NO journal name, GET title.
Abstract,
else journal name,
else title
AeJeT
FOREACH article, GET abstract;
IF NO abstract, GET journal name;
IF NO journal name, GET title.
Journal name,
else abstract,
else title
JeAeT
FOREACH article, GET journal name;
IF NO journal name, GET abstract;
IF NO abstract, GET title.
Abstract & journal name,
else title
AJeT
FOREACH article, GET (abstract, journal name);
IF (NO abstract OR NO journal name), GET title.
Title & abstract,
else title & journal name,
else title
TAeTJeT
FOREACH article, GET (title, abstract);
IF NO abstract, GET (title, journal name);
IF NO journal name, GET title.
Title & journal name,
else title & abstract,
else title
TJeTAeT
FOREACH article, GET (title, journal name);
IF NO journal name, GET (title, abstract);
IF NO abstract, GET title.
Greedy Greedy
FOREACH article, GET(title, abstract, journal name);
IF NO journal name, GET (title, abstract);
IF NO abstract, GET (title, journal name);
IF (NO abstract & NO journal name), GET title.
Table 2: Description of the different corpus types
Appendix C Processing times

### v4

@@ 4.2 Processing Times
Figure LABEL:fig:speed shows the normalized processing times for the different algorithms and corpus types; a table showing the absolute processing times recorded for all classification procedures implemented can be found in appendix B. To improve plot readability, vertical line segments connecting all processing times obtained by each algorithm for the various corpora have been added. Fastest and slowest processing times for each corpus type are shown in seconds at both extremities of the x-axis. Finally, in order to better highlight the differences between both slower and faster algorithms, the x-axis has been rescaled using arcsine transformation.
@@
In the context of the present project, such differences in performance makes processing speed a crucial dimension of linguistic classification performance. Indeed, while a difference of a few seconds at the sample level may not seem so crucial at first glance, the sheer size of OpenAlex implies that processing speed differences beyond one order of magnitude might likely result in delays ranging from several days to a few weeks. Given the fact that the OpenAlex database is continuously updated with new publication data, such delays can cause serious feasability issues.
Regarding the performances of FastSpell and FastText in particular, the former is slightly slower than the latter in all cases, which can be expected given that FastSpell adds an upstream spelling checking functionality to FastText. As shown in the previous figure, this enhancement translates into a gain in recall performance for languages included in the Other category. The question remains, however, whether these local gains in recall performance are worth the resulting costs in both precision and speed.
As for DetectLanguage, the fact that the algorithm can only be used through API queries has an undeniable impact on the processing speed of all procedures using this algorithm: even though the algorithm’s batch mode helps reduce the number of queries by processing multiple inputs at a time, the fact that its processing speed ultimately depends on factors that have nothing to do with algorithmic performance /textitstricto sensu, for example connection speed or server response time, unfortunately makes it an unviable LI strategy for the OpenAlex database.
To a lesser extent, differences in processing times can also be observed between the various corpus types. Across all algorithms, the Titles (T) and Titles & Abstracts (A) corpora are the corpus types whose processing were the fastest and slowest respectively, with times of $15.98$ and $20.64$ seconds. Average document size can probably and partly explain such differences between corpus types: since titles tend to be substantially shorter than abstracts, with $72.26$ characters on average to $1238.85$ characters for the latter, corpora that include abstracts will take more time to process than those that do not. From a performance evaluation perspective, a situation similar the one observed for FastSpell and FastText thus arises: while abstracts can be very informative from a LI perspective, taking them into account for that purpose is however bound to slow down algorithms; whether this gain in information results in an overall gain of performance will ultimately depend on its impact on precision and recall scores, which once again proves the crucial relevance of corpus processing times to the present LI algorithm assessment.
@@ 4.3 Simulation Results
At first glance, what stands out from all the subfigures plotted is the absence of the Titles & Journal Names (J) corpus. Given the visualization method adopted here, this means that no procedure based on that corpus obtained the highest MAP estimate for any algorithm or weighing regime. This contrasts sharply with the results shown in Fig LABEL:fig:perf, which shows that many procedures based on that corpus perform relatively well for certain languages. In our opinion, such performance shift proves that procedures based on other corpus types have shown greater resilience to uncertainty, as modeled by the simulation process.
Looking then at the topmost subfigure, which shows $F_{\beta,\gamma}$ scores for $\gamma=0$ (where is not taken into account), best-performing procedures for each algorithm show interesting similarities. First and foremost, performance for all algorithms seem to follow a dual performance regime, with a transition point at around $\beta\approx 1$, in other words where the weights of precision and recall converge. In the case of DetectLanguage and LangID, this cut-off point occurs a bit earlier ($\beta\approx 0.6$) and later ($\beta\approx 1.2$) along the $\beta$ continuum, but the regime shift is just as sudden and lasting as elsewhere. The first characteristic of that shift is that that all algorithms switch corpus types: whereas the best-performing procedure of CLD3 transitions from the Titles & Abstracts to the Greedy corpus (i.e. it adds journal names), procedures for all other algorithms become title-based as recall weight increases. The second salient feature of the dual-regime observed at $\gamma=0$ is that, again with the exception of the exponential decay of CLD3, which seems unaffected by recall weight increase, all best-performing procedures switch to an exponentially-decreasing growth regime as they hit their respective cut-off points and transitions to the titles-based corpus.
@@
In light of these considerations and the simulation results from which they are derived, which classification procedure performs best depends on the importance given to the different evaluation measures implemented. As regards to corpus types, emphasizing precision over recall will favor the more metadata-rich procedures (i.e. the greedy and titles & abstracts corpora); however, as recall gains in importance and as soon as processing speed is taken into account, the Titles corpus becomes the unanimous choice. As regards to algorithms, while LangID stands out in precision-oriented, speed-overlooking settings, FastText outperforms other algorithms in all other contexts, and the extent of its predominance increases proportionally with both $\beta$ and $\gamma$.
@@ 5 Discussion
The present project aimed to systematically compare various language classification procedures, procedures combining various Python LI algorithms and metadata-based corpora extracted from manually-annotated articles sampled from the OpenAlex database. Following an analysis of precision and recall performance for each algorithm, corpus, and language as well as of processing speeds recorded for each algorithm and corpus type, overall procedure performance at the database level was simulated using probabilistic confusion matrices for each algorithm, corpus, and language as well as a probabilistic model of relative article language frequencies for the whole OpenAlex database. Results show that procedure performance strongly depends on the importance given to each of the measures implemented: for contexts where precision is preferred, using the LangID algorithm on the greedy corpus gives the best results; however, for all cases where recall is considered at least slightly more important than precision or as soon as processing times are given any kind of consideration, the procedure combining the FastText algorithm and the Titles corpus outperforms all other alternatives.
@@
In addition to the quality of its overall performance, both LangID and FastText have the advantage of returning as output the probability value of the most probable language. In this sense, a two- or multi-step classification procedure would be quite feasible. For instance, FastText could be run on the titles of all articles indexed in the database, then other procedures that perform better in specific contexts could be run on all cases where the probability/confidence score of the detected language is below a certain threshold. Given the imbalanced nature of the OpenAlex database and the inherent algorithmic tendency to overpredict the majority class, a special effort could then be made to improve prediction for English documents: as mentioned in cespedes2025evaluating, English documents are overrepresented in the OpenAlex database at a rate of around 9%, which negatively affects the representativity of all other languages in the database to a greater or lesser extent, but also reinforces their underrepresentation in the scholarly and knowledge ecosystem at large. In that perspective and given OpenAlex’s status as the first truly multilingual database, improving the database’s language metadata has social and political implications that go well beyond the functionalities that are normally expected of bibliographic databases.
@@
Results presented and discussed in the previous paragraphs are far from definitive. For once, they are inextricably tied to the current state-of-the-art in the field of automatic LI. The conclusions drawn are in no way meant to solve the problem of the linguistic classification of OpenAlex articles once and for all, but only to attempt to offer the best possible algorithmic recommendation based on current domain knowledge as well as on the available techniques deemed most suitable for that specific purpose.
@@
Another important point to mention is that since the observations and analyses presented here are the direct result of the methodological choices adopted by the authors, the conclusions derived from the results of this project are valid only within the framework of this precise modus operandi. In other words, different choices as regards to the selection of performance measures or the design and parametrization of the simulation process may well have led to different conclusions. For example, the case of the aggregation of performance scores performed here is suggestive in this regard, as increased differences in weighing between precision, recall, and processing speed could lead to different results; the question remains however as to whether more diversified or extreme weighing regimes than those presented here would be really useful. Nevertheless, a certain epistemic caution is in order here, as the results presented here are far from definitive and should not be considered as such.
@@
As regards to OpenAlex itself, it is also important to note that a significant part of the data contained in the database is provided by the journal publishers themselves based on what they consider relevant bibliometric data. Such decisional idiosyncrasies can have an impact on the performance of monolingual LI algorithms and the language classification they provide, for example if multilingual titles or abstracts are provided for specific journals or articles. This specific scenario in turn points to an important consideration regarding the functionalities of the algorithms considered here, but also those of LI algorithms in general. While current monolingual LI algorithms offer above human-level performance over hundreds of different languages, a current and major challenge in Natural Language Processing consists in designing algorithms that achieve high accuracy when dealing with documents written in multiple languages. Given these considerations, multilingual record entries probably represents the biggest current linguistic classification issue for the OpenAlex database: not only is the quality of linguistic metadata inevitably affected by this problem, but its extent is hard to assess, given the difficulty of identifying which attributes are multilingual and as such likely to cause misclassifications.
@@
Despite these limitations, the results presented here show that, with the right LI procedure, a high-quality linguistic classification of all articles indexed in OpenAlex is nevertheless possible, classification which would not only reinforce the latter’s status as the first truly multilingual, large-scale bibliometric database, but also and more generally foster cross-linguistic bibliometric assessments and analyses of unmatched accuracy and comprehensiveness.
Beyond language metadata considerations, we believe the current research project, both in its aims and results, may provide fruitful insights as regards to one one of the biggests problems currently plaguing NLP research, but also science in general, namely the research-practice gap. Summarized in a few words, this gaps refers to the persistent disconnect between what scientific research demonstrates to be effective and what practitioners implement in real-world contexts [haney2024bridging]. Most often, this gap is defined and manifests itself in temporal terms, more precisely as the delay between the publication of research and its integration into professional practice. Studies in healthcare, education, psychology, organizational science, criminology, and social work have reported multi-year lags in putting evidence into use [greenhalgh2016research, lau2016achieving, ebsco2023, elliott2000bridging]. But the gap between research and practice can also further materialize in the partial, inconsistent, unsustained, unscaled or wrongful application of research results. In the worst case, research remains completely dormant, sometimes due to ignorance, but at other times in full awareness, when factors such as lack of budget, knowledge, trust, training, misaligned incentives or simply inertia or resistance to change are at play [rynes2001research, greenhalgh2016research]. More than a mere communication failure, the research-practice gap is a structural divide between two distinct communities with different goals, values, languages, incentives and constraints [Caplan1979_two_communities]. Across domains.and areas of investigations, any such discrepancy has the potential to create a significant divide between research and practice communities or to worsen an already existing one. Ironically, even implementation science itself is facing implementation issues [westerlund2019researchpractice].
As regards to the fields of Artificial Intelligence and Machine Learning, multiple large practitioner surveys by Kaggle, Rexer Analytics, and JetBrains report prevalent, widespread, and enduring usage of regression methods (linear, logistic or generalized linear models) among data scientists in the field [rexer2023, Rexer2020_DataScienceSurvey, Kaggle2020_StateOfML, jetbrains2023]. Thus, while academic ML research produces ever more sophisticated algorithms (deep nets, transformers, etc.), industry practitioners frequently default to simpler, established, reliable, and interpretable methods that favor robustness, simplicity, transparency, explainability, and lower deployment costs.
This structural divide is especially salient in the field of Natural Language Processing. Despite the growing needs for language technology across domains, NLP systems that perform well in research settings often fail to fully translate into practical, sustainable, and scalable deployments [gessler-etal-2025-understanding]. Perhaps here more than elsewhere, this deadlock persists because NLP research seeks to optimize conditions that rarely resemble those encountered at the deployment stage, while practical environments are met with challenges and needs which the research ecosystem and its reward system seldom prioritizes nor even considers.
More than a simple problem of communication, dissemination, or access to knowledge, such misalignment is on the contrary present at all stages of the software development lifecycle, from the earliest stages of a project to its deployment and long-term maintenance. At the requirements stage, research typically defines abstract tasks and goals based on gaps or current trends in their field of expertise; at the other end, healthcare and government adoption studies show that practitioners must satisfy organizational constraints, stakeholder needs, regulatory requirements, and domain-specific risks which are most often completely unrelated with the current state of literature in the relevant research areas [fu2023recommended, jiang2023adoption]. This divergence widens at the data collection and preprocessing: where research builds on static, well-curated benchmarks, practical NLP work must handle messy, heterogeneous, multilingual, and continuously shifting data environments. This “data readiness” gap—long noted by practitioners—means that much of the effort in real-world NLP is spent on data acquisition, cleaning, annotation, documentation, and integration, tasks that the research incentive structure largely overlooks [olsson-sahlgren-2021-datareadiness, joshi2020]. By the time systems reach the modeling and evaluation phase, the mismatch between research and practice is amplified by evaluation regimes that privilege confusion matrix-based performance metrics over robustness, safety, interpretability, and domain fidelity [ribeiro2020checklist, rudin2019stop, jia2017adversarial, bender2021parrots]. As with other research areas, reproducibility considerations are also often overlooked at the evaluation stage: in a compilation of repeatability (reproduction under the same conditions) assessment studies of NLP models , belz2021systematic found out that only 14.03% of all 513 score reproduction attempts were successful, and when reproduction failed, reproduced results were most of the time worse (59.2%) than in the original study.
The various activities involved at the deployment and maintenance phases also contribute to widen the gap. Prototyping rarely accounts for the engineering and infrastructural issues that might arise when models meet production constraints, such as monitoring, logging, versioning, rollback or performance optimization. Especially for large neural models, other practical considerations such as spatial and temporal complexity, latency, scalability, energy consumption, and security risks can create additional and maybe insurmountable barriers [strubell-etal-2019-energy]. In addition, NLP models and ML in general are far more fragile and difficult to maintain than traditional software, as their unstable data dependencies, brittle pipelines, and tightly coupled components make even small changes risky and costly [sculley2015hidden, breck2017mltestscore]. Finally, once NLP systems reach users, new divergences emerge: while research often assumes that the model’s output is the final product, practitioners must design workflows, interfaces, tests, and feedback mechanisms that meet real human needs and allow for safe and iterative improvement [amershi2019guidelines].
Viewed from a software development perspective, the research–practice gap in NLP thus emerges not from a single failure point but from a lifecycle-long accumulation of mismatches in goals, assumptions, incentives, and constraints. Regardless of model performance considerations, each new project that follows this dynamic simply adds an interaction to this vicious circle and widens the gap between researchers and practitioners. To bridge this divide, NLP research has to strive for more than the usual productivity, originality, and accuracy imperatives; as this project tried to show and achieve, it must also and from the outset be geared towards real-world use. This can only be achieved by engaging early and continuously with end-users, organizations, and policy systems throughout the development process and through initiatives like code sharing, co-design and deployment studies [olsson-sahlgren-2021-datareadiness, lin2022truthfulqa, pillai2023toward, gessler-etal-2025-understanding]. Participatory and community-based approaches, which are massively underrepresented in mainstream NLP research [pillai2023toward], can also be very effective in ensuring that research initiatives are both sustainable and useful to the users they target. Contrary to what current incentive structures might suggest, the publication of research results is not an end in itself, but rather the medium through which researchers and practitioners can collaborate to make research truly impactful and innovative. This project, by attempting to provide a comprehensive assessment of automatic language detection algorithms based on the current structure and needs of OpenAlex, aims to contribute to this realignment of NLP research with the objectives and needs of practitioners, likely to reshape research agendas and priorities towards more concretely transformative and truly meliorative outputs.
[…]
Acknowledgements
We would like to thank all of our colleagues who contributed to the annotation of the article sample as part of the article by cespedes2025evaluating. This article stems from the reflections and experiences they shared throughout the annotation process. We also thank Juan Pablo Alperin for his suggestion regarding the use of multi-step language identification procedures.
Declarations
  • Lucia Céspedes, Diego Kozlowski, and Vincent Larivière acknowledge funding from the Social Science and Humanities Research Council of Canada Pan-Canadian Knowledge Access Initiative Grant (Grant Number 1007-2023-0001), and the Fonds de recherche du Québec—Société et Culture through the Programme d’appui aux Chaires UNESCO (Grant Number 338828)
  • The authors declare no conflicts of interest.
  • The data generated and analyzed in this study can either be found in the appendices or in the article by cespedes2025evaluating
  • The code used in this study is available from the corresponding author upon request.
  • CRediT authorship contribution statement. Maxime Holmberg Sainte-Marie: Conceptualization, Data Curation, Formal Analysis, Investigation, Method- ology, Project Administration, Software, Validation, Visualization, Writing - Original Draft; Diego Kozlowski: Conceptualization, Formal Analysis, Methodology; Lucía Céspedes: Conceptualization, Project Administration. Validation, Visualization; Vincent Larivière: Conceptualization, Project Administration, Funding Acquisition, Re- sources, Supervision.
@@ Appendix A OpenAlex linguistic distribution, grouped by metadata availability
Language Title Title & Abstract Title & Journal Name Title, Abstract & Journal Name
@@
Appendix B Processing times